OpenStudy (anonymous):

What is the focus of the parabola given by the equation y = x2 − 2x − 3? Will give medal

OpenStudy (anonymous):

Line is y = -1/2x + 4 Area is 7/3 You need to find the slope of the tangent line first. That is: y' = 2x - 2 = 2(2) - 2 = 2 Perpendicular slope is then -1/2 y - 3 = -1/2(x - 2) y = -1/2x + 4

OpenStudy (anonymous):

Next you integrate, using the line as the upper bound, subtracting the lower bound parabola between 0 and 2 (first quadrant) int(-1/2x + 4 - (x^2 - 2x + 3)) = -1/4x^2 + 4x -x^3/3 + x^2 - 3x Plug in 2 for x and you get -1 + 8 - 8/3 + 4 - 6 12 - 29/3 = 7/3

OpenStudy (anonymous):

@yumfumyum thank you

OpenStudy (anonymous):

your welcome, can you help me

OpenStudy (anonymous):

will give fan and medal

OpenStudy (anonymous):

sure Ill try

OpenStudy (anonymous):

i tagged you

OpenStudy (mathmate):

@judyd94 If you are looking for the focus of the parabola, the answer should be an ordered pair in the form of (x,y). For a parabola \(y=a(x-h)^2+k\) the focus is located at F(h, k+1/(4a) ) You can express \(y=x^2-2x-3\) in the form \(y=a(x-h)^2+k\) by completing the square.

OpenStudy (anonymous):

@mathmate would it be 1,5/14

OpenStudy (mathmate):

1 is correct, since it is the axis of the parabola. the other is not. What do you have for k and a?

OpenStudy (anonymous):

@mathmate I re worked it and got 1,-4